Amgen, Inc., reports the following footnote to its 10-K report. Immunex acquisition. On July 15, 2002, the Company acquired all of the outstanding common stock of Immunex in a transaction accounted for as a business combination. Immunex was a leading biotechnology company dedicated to developing immune system science to protect human health. The acquisition enhanced Amgen’s strategic position within the biotechnology industry by strengthening and diversifying its (1) product base and product pipeline in key therapeutic areas, and (2) discovery research capabilities in proteins and antibodies. The purchase price was allocated to the tangible and identifiable intangible assets acquired and liabilities assumed based on their estimated fair values at the acquisition date. The following table summarizes the estimated fair values of the assets acquired and liabilities assumed as of the acquisition date (in millions):

current assets principally cash and marketable secruitites 1,619.1

deferred tax assets 200.2

property plant and equipment 571.6

inprocess reserach and devlopment 2991.8

indetifiable intangible assets 4803.2

other assets 26.2

current liabilities (579.0)

deferred liabilities (1635.5)

net assets 17,771.8

The allocation of the purchase price was based, in part, on a third-party valuation of the fair values of in-process research and development, identifiable intangible assets, and certain property, plant, and equipment. The estimated fair value of the in-process R&D projects was determined based on the use of a discounted cash flow model. For each project, the estimated after-tax cash flows were probability weighted to take into account the stage of completion and the risks surrounding the successful development and commercialization. These cash flows were then discounted to a present value using discount rates ranging from 12% to 14%.

Of the total assets acquired, what portion is allocated to tangible assets and what portion to intangible assets?

a. 68.6% tangible 31.4% intangible b. 27.2% tangible 72.8% intangible c. 12.1% tangible 87.9% intangible d. 61.2% tangible 38.8% intangible
Answer #1

Answer is option C

c. 12.1% tangible 87.9% intangible

Goodwill = net assets + liabilities – assets = 17771.8+579+1635.5-26.2-4803.2-2991.8-571.6-200.2-1619.1 = 9774.2

Current assets, principally cash and marketable securities

1619.1

Deferred tax assets

200.2

Property, plant, and equipment

571.6

Other assets

26.2

Total tangible assets

2417.1

In-process research and development

2991.8

Identifiable intangible assets, principally developed product technology and core technology

4803.2

Goodwill

9774.2

Total intangible assets

17569.2

Total assets

19986.3

% of tangible assets = total tangible assets / total assets = 2417.1/19986.3 = 12.1%

% of intangible assets = total intangible assets / total assets = 17569.2/19986.3 = 87.9%